### v1

# Sums related to Euler’s totient function

Abstract.
We obtain an upper bound for the sum $\sum_{n\leq N}(a_{n}/\varphi(a_{n}))^{s}$, where $\varphi$ is Euler’s totient function, $s\in\mathbb{N}$, and $a_{1},\ldots,a_{N}$ are positive integers (not necessarily distinct) with some restrictions. As applications, for any $t>0$, we obtain an upper bound for the number of $n\in[1,N]$ such that $a_{n}/\varphi(a_{n})>t$.
Key words and phrases:
Euler’s totient function, prime numbers.
1. Introduction
Let $\varphi$ denote Euler’s totient function and $\log_{2}s=\max(\ln\ln s,1)$. Our main result is the following.
Theorem 1.1.
Let $g$ be the positive multiplicative function such that $g(p)\geq c_{0}>$ for any prime $p$ and
$L=\sum_{p}\frac{1}{pg(p)}<\infty.$
Let $a_{1},\ldots,a_{N}$ be positive integers (not necessarily distinct), $a_{n}\leq M$ for all $1\leq n\leq N$. For $d\in\mathbb{N}$ we set
[…]
Let $\alpha\in(0,1]$, $y=\max((\ln M)^{\alpha},2)$, and $\mathcal{D}$ be the collection of square-free numbers, all of whose prime divisors lie in $(1,y]$ (we note that $1\in\mathcal{D}$). Let
$\omega(n)\leq K\frac{N}{g(n)}$
for any $n\in\mathcal{D}$, where $K>0$ is a constant. Then
$\sum_{n=1}^{N}\bigg{(}\frac{a_{n}}{\varphi(a_{n})}\bigg{)}^{s}\leq\textup{exp}(s\log_{2}s+Cs)N$ (1.1)
for any $s\in\mathbb{N}$. Moreover,
#$\#\big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N.$ (1.2)
Here $C$, $c_{1}$, and $c_{2}$ are positive constants depending only on $K$, $L$, $c_{0}$, and $\alpha$.
From Theorem 1.1 we obtain the following results.
Corollary 1.1.
There are absolute positive constants $C$, $c_{1}$, and $c_{2}$ such that for any $x\geq 2$, $s\in\mathbb{N}$, and $t>0$ we have
$\sum_{p\leq x}\Big{(}\frac{p-1}{\varphi(p-1)}\Big{)}^{s}\leq\textup{exp}(s\log_{2}s+Cs)\pi(x)$
and
#$\#\big{\{}p\leq x:\frac{p-1}{\varphi(p-1)}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))\pi(x).$
Corollary 1.2.
Let $f(n)=b_{d}n^{d}+\ldots+b_{0}$ be a polynomial with integer coefficients such that $b_{d}>0$, $(b_{d},\ldots,b_{0})=1$, and $f:\mathbb{N}\to\mathbb{N}$. Then there exist positive constants $C$, $c_{1}$, and $c_{2}$ depending only on $f$ such that for any $x\geq 1$, $s\in\mathbb{N}$, and $t>0$ we have
$\sum_{n\leq x}\Big{(}\frac{f(n)}{\varphi(f(n))}\Big{)}^{s}\leq\textup{exp}(s\log_{2}s+Cs)x$ (1.3)
and
#$\#\big{\{}n\leq x:\frac{f(n)}{\varphi(f(n))}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))x.$ (1.4)
2. Notation
We reserve the letter $p$ for primes. In particular, the sum $\sum_{p\leq K}$ should be interpreted as being over all prime numbers not exceeding $K$. By $\pi(x)$ we denote the number of primes not exceeding $x$. Let #$\#A$ denote the number of elements of a finite set $A$. By $\mathbb{N}$ we denote the set of all positive integers. Let $(a_{1},\ldots,a_{n})$ be the greatest common divisor of integers $a_{1},\ldots,a_{n}$, and $[a_{1},\ldots,a_{n}]$ the least common multiple of integers $a_{1},\ldots,a_{n}$. For real numbers $x,$ $y$ we also use $(x,y)$ to denote the open interval, and $[x,y]$ to denote the closed interval. The usage of the notation will be clear from the context.
Let $\varphi$ denote Euler’s totient function, i. e.
$\varphi(n)=\#\{1\leq m\leq n:(m,n)=1\},\quad n\in\mathbb{N}.$
We write $\nu(n)$ for the number of distinct primes dividing $n$. By $\mu(n)$ we denote the Möbius mu function, which is defined to be $\mu(n)=(-1)^{\nu(n)}$ if $n$ is square-free, $\mu(n)=0$ otherwise. Let $P^{+}(n)$ denote the greatest prime factor of $n$ (by convention $P^{+}(1)=1$).
By definition, we put
$\sum_{\varnothing}=0,\qquad\prod_{\varnothing}=1.$
The symbol $b|a$ means that $b$ divides $a$. For fixed $a$ the sum $\sum_{b|a}$ and the product $\prod_{b|a}$ should be interpreted as being over all positive divisors of $a$. If $x$ is a real number, then $\lfloor x\rfloor$ denotes its integral part.
3. Proof of Theorem 1.1 and Corollaries 1.1 - LABEL:C3
Proof of Theorem 1.1..
We need the following result.
Lemma 3.1.
There is an absolute positive constant $c$ such that if $\alpha\in(0,1]$ and $n$ is a positive integer, then
$\frac{n}{\varphi(n)}\leq\frac{c}{\alpha}\prod_{p|n:\,\,p\leq(\ln n)^{\alpha}}\biggl{(}1+\frac{1}{p}\biggr{)}.$
Proof.
This is [2, Lemma 3.3]. ∎
For $1\leq n\leq N$, by Lemma 3.1 we have
$\frac{a_{n}}{\varphi(a_{n})}$ $\leq\frac{c}{\alpha}\prod_{p|a_{n}:\,\,p\leq(\ln a_{n})^{\alpha}}\biggl{(}1+\frac{1}{p}\biggr{)}$
$\leq\frac{c}{\alpha}\prod_{p|a_{n}:\,\,p\leq y}\biggl{(}1+\frac{1}{p}\biggr{)}=\frac{c}{\alpha}\sum_{\begin{subarray}{c}d|a_{n}:\\ P^{+}(d)\leq y\end{subarray}}\frac{\mu^{2}(d)}{d}.$
Hence
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}$ $\leq\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{n=1}^{N}\sum_{\begin{subarray}{c}d_{1},\ldots,d_{s}|a_{n}:\\ P^{+}(d_{i})\leq y\end{subarray}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}$
$\leq\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{d_{1},\ldots,d_{s}\in\mathcal{D}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}\sum_{\begin{subarray}{c}1\leq n\leq N:\\ d_{1}|a_{n},\ldots,\,d_{s}|a_{n}\end{subarray}}1$
$=\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{d_{1},\ldots,d_{s}\in\mathcal{D}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}\,\omega([d_{1},\ldots,d_{s}]).$
We obtain
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\leq\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{n\in\mathcal{D}}\omega(n)f(n),$
where
$f(n)=\sum_{\begin{subarray}{c}d_{1},\ldots,d_{s}\in\mathcal{D}:\\ [d_{1},\ldots,d_{s}]=n\end{subarray}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}.$
It is easy to see that $f(n)$ is a multiplicative function supported on $\mathcal{D}$ and
$f(p)=\sum_{k=1}^{s}\binom{s}{k}p^{-k}=(1+p^{-1})^{s}-1$
for any prime $p\leq y$.
Since $\omega(n)\leq KN/g(n)$ for any $n\in\mathcal{D}$, we have
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}$ $\leq KN\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{n\in\mathcal{D}}\frac{f(n)}{g(n)}=KN\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{f(p)}{g(p)}\Big{)}$
$\leq KN\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}.$ (3.1)
For any prime $p>s$, by the mean value theorem there is $\xi\in(0,p^{-1})$ such that
$\Big{(}1+\frac{1}{p}\Big{)}^{s}-1=s(1+\xi)^{s-1}\frac{1}{p}<s\Big{(}1+\frac{1}{s}\Big{)}^{s-1}\frac{1}{p}<\frac{es}{p}.$
Since $\ln(1+x)\leq x$ for any $x\geq 0$, we obtain
$\prod_{p>s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\prod_{p>s}\Big{(}1+\frac{es}{pg(p)}\Big{)}\leq\textup{exp}\Big{(}es\sum_{p>s}\frac{1}{pg(p)}\Big{)}\leq\textup{exp}(eLs).$ (3.2)
If $p\leq s$, then
$1+\frac{(1+p^{-1})^{s}-1}{g(p)}<1+\frac{(1+p^{-1})^{s}}{c_{0}}\leq\Big{(}1+\frac{1}{c_{0}}\Big{)}(1+p^{-1})^{s}\leq\Big{(}1+\frac{1}{c_{0}}\Big{)}\textup{exp}\Big{(}\frac{s}{p}\Big{)}.$
By Mertens’ theorem we have
$\prod_{p\leq s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$ $\leq\Big{(}1+\frac{1}{c_{0}}\Big{)}^{s}\textup{exp}\Big{(}\sum_{p\leq s}\frac{s}{p}\Big{)}$
$\leq\Big{(}1+\frac{1}{c_{0}}\Big{)}^{s}\textup{exp}(s\log_{2}s+c_{3}s),$ (3.3)
where $c_{3}>0$ is an absolute constant. From (3.1), (3.2), and (3.3) we obtain
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\leq\textup{exp}(s\log_{2}s+Cs)N,$
where $C$ is a positive constant depending only on $K$, $L$, $c_{0}$, and $\alpha$. The inequality (1.1) is proved.
Now we prove the inequality (1.2). We have
$\textup{exp}(s\log_{2}s+Cs)N$ $\geq\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\geq\sum_{n\leq N:\,a_{n}/\varphi(a_{n})>t}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}$
$\geq t^{s}\sum_{n\leq N:\,a_{n}/\varphi(a_{n})>t}1.$
We obtain
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}\leq\textup{exp}(s\log_{2}s+Cs-s\ln t)N.$
We take
$s=\lfloor\textup{exp}(te^{-(C+1)})\rfloor+1.$
It is easy to see that
$s\log_{2}s+Cs-s\ln t=-\textup{exp}(te^{-(C+1)})+O(\ln t)\leq-\textup{exp}(te^{-(C+1)}/2),$
if $t\geq t_{0}$ (here $t_{0}=t_{0}(K,L,c_{0},\alpha)$ is a positive constant depending only on $K$, $L$, $c_{0}$, and $\alpha$). We set
$c_{2}=e^{-(C+1)}/2,\qquad c_{1}=\max(\textup{exp}(\textup{exp}(c_{2}t_{0})),1).$
Then $c_{1}$ and $c_{2}$ are positive constants depending only on $K$, $L$, $c_{0}$, and $\alpha$. If $t\leq t_{0}$, then
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}$ $\leq N\leq N\,\frac{\textup{exp}(\textup{exp}(c_{2}t_{0}))}{\textup{exp}(\textup{exp}(c_{2}t))}$
$\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N.$
We obtain
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N$
for any $t>0$, and the inequality (1.2) is proved. This completes the proof of Theorem 1.1. ∎
Proof of Corollary 1.1..
We set $a_{n}=p_{n}-1$, where $p_{n}$ is the $n$th prime. It is clear that $N=\pi(x)$. We take $\alpha=1/2$, $M=x$. We can assume that $x\geq x_{0}$, where $x_{0}>0$ is large constant. If $n\in\mathcal{D}$, then
$n\leq\prod_{p\leq y}p=\textup{exp}((1+o(1))y)\leq\textup{exp}(2(\ln x)^{1/2})\leq\sqrt{x}.$
By the Brun - Titchmarsh inequality we have
$\omega(n)=\#\{p\leq x:p\equiv 1\text{ {(}mod $n${)}}\}\leq\frac{3x}{\varphi(n)\ln(2x/n)}\leq K\,\frac{\pi(x)}{\varphi(n)},$
where $K$ is an absolute positive constant. Since $\varphi(n)$ is a positive multiplicative function, $\varphi(p)=p-1\geq 1$ for any prime $p$ (hence, we can take $c_{0}=1$), and
$L=\sum_{p}\frac{1}{p\varphi(p)}=\sum_{p}\frac{1}{p(p-1)}<\infty,$
we can apply Theorem 1.1. Since $K$, $L$, $c_{0}$, and $\alpha$ are absolute positive constants, the constants $C$, $c_{1}$, and $c_{2}$ are absolute and positive. Corollary 1.1 is proved. ∎
Proof of Corollary 1.2..
We set $a_{n}=f(n)$. It is clear that $N=\lfloor x\rfloor$. There exists a positive constant $\gamma$ depending only on $f$ such that $f(n)\leq\gamma n^{d}$ for any positive integer $n$. We take $M=\gamma x^{d}$ and $\alpha=1/2$. We can assume that $x\geq x_{0}$, where $x_{0}>0$ is large constant depending only on $f$. We have
$y=(\ln M)^{\alpha}=(d\ln x+\ln\gamma)^{1/2}\leq(\ln x)^{3/4},$
if $x_{0}$ is large enough. If $n\in\mathcal{D}$, then
$n\leq\prod_{p\leq y}p\leq\textup{exp}(2y)\leq\textup{exp}(2(\ln x)^{3/4})\leq\sqrt{x}<\lfloor x\rfloor=N.$
We need the following result.
Lemma 3.2.
Let $d$ and $m$ be positive integers. Let
$f(x)=\sum_{i=0}^{d}b_{i}x^{i},$
where $b_{0},\ldots,b_{d}$ are integers with $(b_{0},\ldots,b_{d},m)=1$. Let $\rho(f,m)$ denote the number of solutions of the congruence $f(x)\equiv 0$ (mod $m$). Then
$\rho(f,m)\leq cdm^{1-1/d},$
where $c>0$ is an absolute constant.
Proof.
This is [1, Theorem 2]. ∎
By Lemma 3.2 we have
$\omega(n)=\#\{k\leq N$ $:f(k)\equiv 0\text{ {(}mod $n${)}}\}\leq\rho(f,n)\Big{(}\frac{N}{n}+1\Big{)}\leq 2\rho(f,n)\,\frac{N}{n}$
$\leq(2c)dn^{1-1/d}\,\frac{N}{n}=K\,\frac{N}{n^{1/d}},$
where $K=2cd$. The function $g(n)=n^{1/d}$ is positive and multiplicative. Also, $g(p)=p^{1/d}\geq 1$ for any prime $p$ (hence, we can take $c_{0}=1$) and
$L=\sum_{p}\frac{1}{pg(p)}=\sum_{p}\frac{1}{p^{1+1/d}}<\infty.$
By Theorem 1.1 there exist positive constants $C$, $c_{1}$, and $c_{2}$ (depending only on $d$) such that (1.3) and (1.4) hold. This completes the proof of Corollary 1.2.
4. Acknowledgements
The author is grateful to Mikhail R. Gabdullin for useful comments.
This research was supported by Russian Science Foundation, grant 20-11-20203, https://rscf.ru/en/project/20-11-20203/.