### v2

# Sums related to Euler’s totient function

Abstract.
We obtain an upper bound for the sum $\sum_{n\leq N}(a_{n}/\varphi(a_{n}))^{s}$, where $\varphi$ is Euler’s totient function, $s\in\mathbb{N}$, and $a_{1},\ldots,a_{N}$ are positive integers (not necessarily distinct) with some restrictions. As applications, for any $t>0$, we obtain an upper bound for the number of $n\in[1,N]$ such that $a_{n}/\varphi(a_{n})>t$.
Key words and phrases:
Euler’s totient function, multiplicative functions, prime numbers.
1. Introduction
Let $\varphi$ denote Euler’s totient function. We prove
Theorem 1.1.
Let $a_{1},\ldots,a_{N}$ be positive integers (not necessarily distinct), $a_{n}\leq M$ for all $1\leq n\leq N$. For $d\in\mathbb{N}$ we set
[…]
Let $\alpha\in(0,1]$, $y=\max((\ln M)^{\alpha},2)$, and $\mathcal{D}$ be the collection of square-free numbers, all of whose prime divisors lie in $(1,y]$ (we note that $1\in\mathcal{D}$). Let
$\omega(n)\leq\frac{K}{g(n)}$
for any $n\in\mathcal{D}$, where $K>0$ is a constant (depending on $N$) and $g(n)$ is a multiplicative function. Then
$\sum_{n=1}^{N}\bigg{(}\frac{a_{n}}{\varphi(a_{n})}\bigg{)}^{s}\leq K\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$
for any $s\in\mathbb{N}$. Here $c>0$ is an absolute constant.
Theorem 1.2.
Suppose that the assumptions of Theorem 1.1 hold, $K=\gamma N$, $g(p)\geq c_{0}>0$ for any prime $p$, and
$L=\sum_{p}\frac{1}{pg(p)}<\infty.$
Then there are positive constants $C$, $c_{1}$, and $c_{2}$ depending only on $\gamma$, $L$, $c_{0}$, and $\alpha$ such that
$\sum_{n=1}^{N}\bigg{(}\frac{a_{n}}{\varphi(a_{n})}\bigg{)}^{s}\leq\textup{exp}(s\ln\ln(s+2)+Cs)N$ (1.1)
for any $s\in\mathbb{N}$, and
#$\#\big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N$ (1.2)
for any $t>0$.
From Theorems 1.1 and 1.2 we obtain the following results.
Corollary 1.1.
Let $f(n)=b_{d}n^{d}+\ldots+b_{0}$ be a polynomial with integer coefficients such that $b_{d}>0$, $(b_{d},\ldots,b_{0})=1$, and $f:\mathbb{N}\to\mathbb{N}$. Then there exist positive constants $C$, $c_{1}$, and $c_{2}$ depending only on $f$ such that for any $x\geq 1$, $s\in\mathbb{N}$, and $t>0$ we have
$\sum_{n\leq x}\Big{(}\frac{f(n)}{\varphi(f(n))}\Big{)}^{s}\leq\textup{exp}(s\ln\ln(s+2)+Cs)x$ (1.3)
and
#$\#\big{\{}n\leq x:\frac{f(n)}{\varphi(f(n))}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))x.$ (1.4)
Corollary 1.1 extends Corollary 1.1 in [3] which showed the inequality (1.3) but with an upper bound $\textup{exp}(s\ln s+Cs)x$ and without the inequality (1.4).
Let =$\mathcal{L}=\{L_{1},\ldots,L_{k}\}$ be a set of $k$ linear functions with integer coefficients
$L_{i}(n)=a_{i}n+b_{i},\qquad i=1,\ldots,k.$
For $L(n)=an+b$, $a,b\in\mathbb{Z}$, we define
$\Delta_{L}=|a|\prod_{i=1}^{k}|ab_{i}-ba_{i}|.$
We note that if $a_{i}=a$ for all $1\leq i\leq k$, then
$\Delta_{L}=|a|^{k+1}\prod_{i=1}^{k}|b-b_{i}|.$
Modern application of the sieve methods involves the sums
$\sum_{(a,b)\in\Omega}\frac{\Delta_{L}}{\varphi(\Delta_{L})}$
(see, for example, [2]). Here $(a,b)$ denotes a vector and $\Omega$ is a finite set in $\mathbb{Z}^{2}$. We prove
Corollary 1.2.
Let $s$, $a$, $k\geq 2$, $b_{1}<\ldots<b_{k}$ be positive integers. Let $x\geq 3$, $(\ln x)^{-9/10}\leq\eta\leq 1$, and $|b_{i}|\leq\ln x$ for all $1\leq i\leq k$. We set
$f(b)=(b-b_{1})\dots(b-b_{k}).$
Then the following statements hold.
1) If $k\geq\ln\ln x$ or $s\leq k$, then
$\sum_{\begin{subarray}{c}|b|\leq\eta\ln x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}c\,\frac{a}{\varphi(a)}\ln k\Big{)}^{s}\eta\ln x.$ (1.5)
2) If $2\leq k<\ln\ln x$ and $s>k$, then
$\sum_{\begin{subarray}{c}|b|\leq\eta\ln x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}c\,\frac{a}{\varphi(a)}\ln s\Big{)}^{s}\eta\ln x.$ (1.6)
Here $c>0$ is an absolute constant.
Corollary 1.2 extends a result of Maynard ([2, Lemma 8.1]) which showed the same result but with $s=1$ and extends Theorem 1.4 in [3] which showed the inequality (1.6) but with an upper bound
$\Big{(}c\,\frac{a}{\varphi(a)}\ln k\Big{)}^{s}s!\,\eta\ln x.$
From Theorem 1.2 we obtain
Corollary 1.3.
There are absolute positive constants $C$, $c_{1}$, and $c_{2}$ such that for any $x\geq 2$, $s\in\mathbb{N}$, and $t>0$ we have
$\sum_{p\leq x}\Big{(}\frac{p-1}{\varphi(p-1)}\Big{)}^{s}\leq\textup{exp}(s\ln\ln(s+2)+Cs)\pi(x)$ (1.7)
and
#$\#\big{\{}p\leq x:\frac{p-1}{\varphi(p-1)}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))\pi(x).$ (1.8)
2. Notation
We reserve the letter $p$ for primes. In particular, the sum $\sum_{p\leq K}$ should be interpreted as being over all prime numbers not exceeding $K$. By $\pi(x)$ we denote the number of primes not exceeding $x$. Let #$\#A$ denote the number of elements of a finite set $A$. By $\mathbb{N}$ we denote the set of all positive integers. Let $(a_{1},\ldots,a_{n})$ be the greatest common divisor of integers $a_{1},\ldots,a_{n}$, and $[a_{1},\ldots,a_{n}]$ the least common multiple of integers $a_{1},\ldots,a_{n}$. For real numbers $x,$ $y$ we also use $(x,y)$ to denote the open interval, and $[x,y]$ to denote the closed interval. The usage of the notation will be clear from the context.
Let $\varphi$ denote Euler’s totient function, i. e.
$\varphi(n)=\#\{1\leq m\leq n:(m,n)=1\},\quad n\in\mathbb{N}.$
We write $\nu(n)$ for the number of distinct primes dividing $n$. By $\mu(n)$ we denote the Möbius mu function, which is defined to be $\mu(n)=(-1)^{\nu(n)}$ if $n$ is square-free, $\mu(n)=0$ otherwise. Let $P^{+}(n)$ denote the greatest prime factor of $n$ (by convention $P^{+}(1)=1$).
By definition, we put
$\sum_{\varnothing}=0,\qquad\prod_{\varnothing}=1.$
The symbol $b|a$ means that $b$ divides $a$. For fixed $a$ the sum $\sum_{b|a}$ and the product $\prod_{b|a}$ should be interpreted as being over all positive divisors of $a$. If $x$ is a real number, then $[x]$ denotes its integral part and $\{x\}=x-[x]$.
3. Proof of Theorems 1.1 and 1.2
Proof of Theorem 1.1..
We need the following result.
Lemma 3.1.
There is an absolute positive constant $c$ such that if $\alpha\in(0,1]$ and $n$ is a positive integer, then
$\frac{n}{\varphi(n)}\leq\frac{c}{\alpha}\prod_{p|n:\,\,p\leq(\ln n)^{\alpha}}\biggl{(}1+\frac{1}{p}\biggr{)}.$
Proof.
This is [3, Lemma 3.3]. ∎
For $1\leq n\leq N$, by Lemma 3.1 we have
$\frac{a_{n}}{\varphi(a_{n})}$ $\leq\frac{c}{\alpha}\prod_{p|a_{n}:\,\,p\leq(\ln a_{n})^{\alpha}}\biggl{(}1+\frac{1}{p}\biggr{)}$
$\leq\frac{c}{\alpha}\prod_{p|a_{n}:\,\,p\leq y}\biggl{(}1+\frac{1}{p}\biggr{)}=\frac{c}{\alpha}\sum_{\begin{subarray}{c}d|a_{n}:\\ P^{+}(d)\leq y\end{subarray}}\frac{\mu^{2}(d)}{d}.$
Hence
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}$ $\leq\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{n=1}^{N}\sum_{\begin{subarray}{c}d_{1},\ldots,d_{s}|a_{n}:\\ P^{+}(d_{i})\leq y\end{subarray}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}$
$\leq\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{d_{1},\ldots,d_{s}\in\mathcal{D}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}\sum_{\begin{subarray}{c}1\leq n\leq N:\\ d_{1}|a_{n},\ldots,\,d_{s}|a_{n}\end{subarray}}1$
$=\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{d_{1},\ldots,d_{s}\in\mathcal{D}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}\,\omega([d_{1},\ldots,d_{s}]).$
We obtain
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\leq\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{n\in\mathcal{D}}\omega(n)f(n),$
where
$f(n)=\sum_{\begin{subarray}{c}d_{1},\ldots,d_{s}\in\mathcal{D}:\\ [d_{1},\ldots,d_{s}]=n\end{subarray}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}.$
It is easy to see that $f(n)$ is supported on $\mathcal{D}$. Suppose that $m,n\in\mathcal{D}$ and $(m,n)=1$. For any $d_{1},\ldots,d_{s}\in\mathcal{D}$ such that $[d_{1},\ldots,d_{s}]=mn$ there are unique $d_{i}^{{}^{\prime}},d_{i}^{{}^{\prime\prime}}\in\mathcal{D},$ $i=1,\ldots,s,$ such that $d_{i}^{{}^{\prime}}d_{i}^{{}^{\prime\prime}}=d_{i}$, $(d_{i}^{{}^{\prime}},d_{i}^{{}^{\prime\prime}})=1$, $[d_{1}^{{}^{\prime}},\ldots,d_{s}^{{}^{\prime}}]=n$, and $[d_{1}^{{}^{\prime\prime}},\ldots,d_{s}^{{}^{\prime\prime}}]=m$. Hence
$f(nm)=\sum_{\begin{subarray}{c}d_{1}^{{}^{\prime}},\ldots,d_{s}^{{}^{\prime}}\in\mathcal{D}:\\ [d_{1}^{{}^{\prime}},\ldots,d_{s}^{{}^{\prime}}]=n\end{subarray}}\frac{\mu^{2}(d_{1}^{{}^{\prime}})\ldots\mu^{2}(d_{s}^{{}^{\prime}})}{d_{1}^{{}^{\prime}}\ldots d_{s}^{{}^{\prime}}}\sum_{\begin{subarray}{c}d_{1}^{{}^{\prime\prime}},\ldots,d_{s}^{{}^{\prime\prime}}\in\mathcal{D}:\\ [d_{1}^{{}^{\prime\prime}},\ldots,d_{s}^{{}^{\prime\prime}}]=m\end{subarray}}\frac{\mu^{2}(d_{1}^{{}^{\prime\prime}})\ldots\mu^{2}(d_{s}^{{}^{\prime\prime}})}{d_{1}^{{}^{\prime\prime}}\ldots d_{s}^{{}^{\prime\prime}}}=f(n)f(m).$
Since $f(n)=0$ for any $n\notin\mathcal{D}$, we obtain that $f(n)$ is a multiplicative function. It is clear that
$f(p)=\sum_{k=1}^{s}\binom{s}{k}p^{-k}=(1+p^{-1})^{s}-1$
for any prime $p\leq y$.
Since $\omega(n)\leq K/g(n)$ for any $n\in\mathcal{D}$, we have
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}$ $\leq K\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{n\in\mathcal{D}}\frac{f(n)}{g(n)}=K\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{f(p)}{g(p)}\Big{)}$
$=K\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}.$
Theorem 1.1 is proved. ∎
Proof of Theorem 1.2..
Since $g(p)>0$ for any prime $p$, from Theorem 1.1 we obtain
$\sum_{n=1}^{N}\bigg{(}\frac{a_{n}}{\varphi(a_{n})}\bigg{)}^{s}$ $\leq\gamma N\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$
$\leq\gamma N\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}.$ (3.1)
For any prime $p>s$, by the mean value theorem there is $\xi\in(0,p^{-1})$ such that
$\Big{(}1+\frac{1}{p}\Big{)}^{s}-1=s(1+\xi)^{s-1}\frac{1}{p}<s\Big{(}1+\frac{1}{s}\Big{)}^{s-1}\frac{1}{p}<\frac{es}{p}.$ (3.2)
Since $\ln(1+x)\leq x$ for any $x\geq 0$, we obtain
$\prod_{p>s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\prod_{p>s}\Big{(}1+\frac{es}{pg(p)}\Big{)}\leq\textup{exp}\Big{(}es\sum_{p>s}\frac{1}{pg(p)}\Big{)}\leq\textup{exp}(eLs).$ (3.3)
If $p\leq s$, then
$1+\frac{(1+p^{-1})^{s}-1}{g(p)}<1+\frac{(1+p^{-1})^{s}}{c_{0}}\leq\Big{(}1+\frac{1}{c_{0}}\Big{)}(1+p^{-1})^{s}\leq\Big{(}1+\frac{1}{c_{0}}\Big{)}\textup{exp}\Big{(}\frac{s}{p}\Big{)}.$
By Mertens’ theorem we have
$\prod_{p\leq s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$ $\leq\Big{(}1+\frac{1}{c_{0}}\Big{)}^{s}\textup{exp}\Big{(}\sum_{p\leq s}\frac{s}{p}\Big{)}$
$\leq\Big{(}1+\frac{1}{c_{0}}\Big{)}^{s}\textup{exp}(s\ln\ln(s+2)+c_{3}s),$ (3.4)
where $c_{3}>0$ is an absolute constant. From (3.1), (3.3), and (3.4) we obtain
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\leq\textup{exp}(s\ln\ln(s+2)+Cs)N,$
where $C$ is a positive constant depending only on $\gamma$, $L$, $c_{0}$, and $\alpha$. The inequality (1.1) is proved.
Now we prove the inequality (1.2). We have
$\textup{exp}(s\ln\ln(s+2)+Cs)N$ $\geq\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\geq\sum_{n\leq N:\,a_{n}/\varphi(a_{n})>t}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}$
$\geq t^{s}\sum_{n\leq N:\,a_{n}/\varphi(a_{n})>t}1.$
We obtain
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}\leq\textup{exp}(s\ln\ln(s+2)+Cs-s\ln t)N.$ (3.5)
We take
$s=[\textup{exp}(te^{-(C+1)})]+1.$
Then
$s+2=\textup{exp}(te^{-(C+1)})+3-\theta,\text{ where }\theta:=\{\textup{exp}(te^{-(C+1)})\},$
and
$\ln(s+2)$ $=te^{-(C+1)}+\ln\Big{(}1+\frac{3-\theta}{\textup{exp}(te^{-(C+1)})}\Big{)}$
$=te^{-(C+1)}+R_{1},\qquad 0<R_{1}\leq\frac{3}{\textup{exp}(te^{-(C+1)})}.$
We get
$\ln\ln(s+2)$ $=\ln t-(C+1)+\ln\Big{(}1+\frac{e^{(C+1)}R_{1}}{t}\Big{)}$
$=\ln t-(C+1)+R_{2},\qquad 0<R_{2}\leq\frac{3e^{C+1}}{t\,\textup{exp}(te^{-(C+1)})}.$
Therefore
$s\ln\ln(s+2)+Cs-s\ln t=-s+sR_{2},$
where
$0<sR_{2}\leq\frac{3e^{C+1}(\textup{exp}(te^{-(C+1)})+1)}{t\,\textup{exp}(te^{-(C+1)})}\leq\frac{6e^{C+1}}{t}.$
We obtain
$s\ln\ln(s+2)+Cs-s\ln t\leq-\textup{exp}(te^{-(C+1)})+\frac{6e^{C+1}}{t}\leq-\textup{exp}(te^{-(C+1)}/2),$ (3.6)
if $t\geq t_{0}$ (here $t_{0}=t_{0}(C)=t_{0}(\gamma,L,c_{0},\alpha)$ is a positive constant depending only on $\gamma$, $L$, $c_{0}$, and $\alpha$).
We set
$c_{2}=e^{-(C+1)}/2,\qquad c_{1}=\max(\textup{exp}(\textup{exp}(c_{2}t_{0})),1).$
Then $c_{1}$ and $c_{2}$ are positive constants depending only on $\gamma$, $L$, $c_{0}$, and $\alpha$. From (3.5) and (3.6) we obtain
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}\leq\textup{exp}(-\textup{exp}(c_{2}t))N,$
if $t>t_{0}$. If $0<t\leq t_{0}$, then
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}$ $\leq N\leq N\,\frac{\textup{exp}(\textup{exp}(c_{2}t_{0}))}{\textup{exp}(\textup{exp}(c_{2}t))}$
$\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N.$
We obtain
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N$
for any $t>0$, and the inequality (1.2) is proved. This completes the proof of Theorem 1.2. ∎
4. Proof of Corollaries 1.3 - 1.2
Proof of Corollary 1.1..
We set $a_{n}=f(n)$. It is clear that $N=[x]$. There exists a positive constant $\tau$ depending only on $f$ such that $f(n)\leq\tau n^{d}$ for any positive integer $n$. We take $M=\tau x^{d}$ and $\alpha=1/2$. We can assume that $x\geq x_{0}$, where $x_{0}>0$ is a large constant depending only on $f$. We have
$y=(\ln M)^{\alpha}=(d\ln x+\ln\tau)^{1/2}\leq(\ln x)^{3/4},$
if $x_{0}$ is large enough. If $n\in\mathcal{D}$, then
$n\leq\prod_{p\leq y}p\leq\textup{exp}(2y)\leq\textup{exp}(2(\ln x)^{3/4})\leq\sqrt{x}<[x]=N.$
We need the following result.
Lemma 4.1.
Let $d$ and $m$ be positive integers. Let
$f(x)=\sum_{i=0}^{d}b_{i}x^{i},$
where $b_{0},\ldots,b_{d}$ are integers with $(b_{0},\ldots,b_{d},m)=1$. Let $\rho(f,m)$ denote the number of solutions of the congruence $f(x)\equiv 0$ (mod $m$). Then
$\rho(f,m)\leq cdm^{1-1/d},$
where $c>0$ is an absolute constant.
Proof.
This is [1, Theorem 2]. ∎
By Lemma 4.1 we have
$\omega(n)=\#\{k\leq N$ $:f(k)\equiv 0\text{ {(}mod $n${)}}\}\leq\rho(f,n)\Big{(}\frac{N}{n}+1\Big{)}\leq 2\rho(f,n)\,\frac{N}{n}$
$\leq(2c)dn^{1-1/d}\,\frac{N}{n}=\frac{\gamma N}{n^{1/d}},$
where $\gamma=2cd$. The function $g(n)=n^{1/d}$ is multiplicative. Also, $g(p)=p^{1/d}\geq 1$ for any prime $p$ (hence, we can take $c_{0}=1$) and
$L=\sum_{p}\frac{1}{pg(p)}=\sum_{p}\frac{1}{p^{1+1/d}}<\infty.$
By Theorem 1.2 there exist positive constants $C$, $c_{1}$, and $c_{2}$ (depending only on $d$) such that (1.3) and (1.4) hold. This completes the proof of Corollary 1.1.
Proof of Corollary 1.2..
We can assume that $x\geq x_{0}$, where $x_{0}$ is a large absolute constant. We set
$\Omega=\{b\in\mathbb{Z}:|b|\leq\eta\ln x,b\neq b_{1},\ldots,b_{k}\}.$
Since $\varphi(mn)\geq\varphi(m)\varphi(n)$ for all positive integers $m$ and $n$, and $a^{k+1}/\varphi(a^{k+1})=a/\varphi(a)$, we obtain
$S=\sum_{b\in\Omega}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}\frac{a}{\varphi(a)}\Big{)}^{s}\sum_{b\in\Omega}\Big{(}\frac{|f(b)|}{\varphi(|f(b)|)}\Big{)}^{s}=\Big{(}\frac{a}{\varphi(a)}\Big{)}^{s}S_{0}.$ (4.1)
Let $b\in\Omega$. Hence, $|b-b_{i}|\leq 2\ln x$ for all $1\leq i\leq k$, and $|f(b)|\leq(2\ln x)^{k}$. Suppose that $k\geq\ln\ln x$. Since $n/\varphi(n)\leq c\ln\ln(n+2)$ for any positive integer $n$, where $c$ is an absolute positive constant, we obtain
$\frac{|f(b)|}{\varphi(|f(b)|)}\leq c(\ln k+\ln\ln(4\ln x))\leq c_{1}\ln k.$
Hence,
$S\leq\Big{(}\frac{a}{\varphi(a)}\,c_{1}\ln k\Big{)}^{s}\#\Omega\leq\Big{(}c_{2}\,\frac{a}{\varphi(a)}\ln k\Big{)}^{s}\eta\ln x,$
where $c_{2}$ is an absolute positive constant.
Suppose that $2\leq k<\ln\ln x$. We take $\alpha=1/4$, $M=(2\ln x)^{k}$. Then
$y=(\ln M)^{\alpha}\leq 2(\ln\ln x)^{1/2}.$
If $n\in\mathcal{D}$, then
$n\leq\prod_{p\leq y}p\leq\textup{exp}(2y)\leq\textup{exp}(4(\ln\ln x)^{1/2})\leq\eta\ln x.$
For $n\in\mathcal{D}$, we have
$\omega(n)=\#\{b\in\Omega:f(b)\equiv 0\text{ (mod $n$)}\}$ $\leq\Big{(}\frac{2\eta\ln x}{n}+1\Big{)}\prod_{p|n}\min(p,k)$
$\leq\frac{3\eta\ln x}{n}\prod_{p|n}\min(p,k).$
We set $g(1)=1,$
$g(p^{\beta})=\begin{cases}p/\min(p,k),&\text{if $\beta=1$;}\\ 1,&\text{if $\beta\geq 2$};\end{cases}$
and $g(p_{1}^{\beta_{1}}\dots p_{r}^{\beta_{r}})=g(p_{1}^{\beta_{1}})\dots g(p_{r}^{\beta_{r}})$, where $p_{1},\ldots,p_{r}$ are pairwise distinct prime numbers. Then $g(n)$ is a multiplicative function and
$\omega(n)\leq\frac{3\eta\ln x}{g(n)}$
for any $n\in\mathcal{D}$. Since $g(p)>0$ for any prime $p$, from Theorem 1.1 we obtain
$\sum_{b\in\Omega}\Big{(}\frac{|f(b)|}{\varphi(|f(b)|)}\Big{)}^{s}\leq c^{s}(\eta\ln x)\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$
$\leq c^{s}(\eta\ln x)\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$ (4.2)
(here $c>0$ is an absolute constant). By (4.1) and (4.2) we have
$S\leq\Big{(}c\,\frac{a}{\varphi(a)}\Big{)}^{s}\eta\ln x\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}.$
Suppose that $s\leq k$ (we recall that $2\leq k<\ln\ln x$). If $p>k$, then $p>s$ and (3.2) holds. Since $g(p)=p/k$, we obtain
$\prod_{p>k}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\prod_{p>k}\Big{(}1+\frac{eks}{p^{2}}\Big{)}\leq c_{1}^{s},$
where $c_{1}>0$ is an absolute constant. If $p\leq k$, then $g(p)=1$ and we have
$1+\frac{(1+p^{-1})^{s}-1}{g(p)}=(1+p^{-1})^{s}\leq\textup{exp}\Big{(}\frac{s}{p}\Big{)}.$
Hence
$\prod_{p\leq k}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\textup{exp}\Big{(}\sum_{p\leq k}\frac{s}{p}\Big{)}\leq(c_{2}\ln k)^{s},$
where $c_{2}>0$ is an absolute constant. The inequality (1.5) is proved.
Suppose that $s>k$ and $2\leq k<\ln\ln x$. Since $g(p)\geq p/k>p/s$ for any prime $p$, we obtain
$\prod_{p>s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\prod_{p>s}\Big{(}1+\frac{es^{2}}{p^{2}}\Big{)}\leq c_{3}^{s}.$
Since $g(p)\geq 1$ for any prime $p$, we have
$\prod_{p\leq s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\textup{exp}\Big{(}\sum_{p\leq s}\frac{s}{p}\Big{)}\leq(c_{4}\ln s)^{s}.$
The inequality (1.6) is proved. This completes the proof of Corollary 1.2.
Proof of Corollary 1.3..
We set $a_{n}=p_{n}-1$, where $p_{n}$ is the $n$th prime. It is clear that $N=\pi(x)$. We take $\alpha=1/2$, $M=x$. We can assume that $x\geq x_{0}$, where $x_{0}$ is a large absolute constant. Then $y=(\ln x)^{1/2}$. If $n\in\mathcal{D}$, then
$n\leq\prod_{p\leq y}p=\textup{exp}((1+o(1))y)\leq\textup{exp}(2(\ln x)^{1/2})\leq\sqrt{x}.$
By the Brun - Titchmarsh inequality (see, for example, [4, Exercise 9.3.13]) we have
$\omega(n)=\#\{p\leq x:p\equiv 1\text{ {(}mod $n${)}}\}\leq\frac{3x}{\varphi(n)\ln(2x/n)}\leq\gamma\,\frac{\pi(x)}{\varphi(n)},$
where $\gamma$ is an absolute positive constant. Since $\varphi(n)$ is a multiplicative function, $\varphi(p)=p-1\geq 1$ for any prime $p$ (hence, we can take $c_{0}=1$), and
$L=\sum_{p}\frac{1}{p\varphi(p)}=\sum_{p}\frac{1}{p(p-1)}<\infty,$
we can apply Theorem 1.2. Since $\gamma$, $L$, $c_{0}$, and $\alpha$ are absolute positive constants, the constants $C$, $c_{1}$, and $c_{2}$ in (1.7) and (1.8) are absolute and positive. Corollary 1.3 is proved. ∎
5. Acknowledgements
The author is grateful to Mikhail R. Gabdullin for useful comments.
This research was supported by Russian Science Foundation, grant 20-11-20203, https://rscf.ru/en/project/20-11-20203/.